Bill is 1 year less than twice as old as his sister Caroline.
The total of their ages is 26.
How old is Bill now? Please help!
Thanks1

Answer by xcentaur(357)

Let Bill be x years old,let Caroline be y years old.
Now,
"Bill is 1 year less than twice as old as his sister Caroline."


twice as old as his sister Caroline means (2y)

Bill is 1 year less than twice as old as his sister means (2y-1)


(2y-1) is Bill's age,which is 'x'
x=2y-1


"The total of their ages is 26"
that means
x+y=26
we know x=2y-1,so put this value in the above equation
x+y=26
(2y-1)+y=26
3y-1=26
3y=26+1
3y=27
y=9
x=2y-1=2(9)-1=18-1=17


Bill:17 years old,Caroline:9 years old


Hope this helps,
Prabhat